What  is the electric field due to a point charge of  45  μC   at a distance of 1 meter away from it?   


1
Expert's answer
2021-04-28T09:55:59-0400

The electric field due to a point charge "q = 45\mu C = 45\times 10^{-6}C" at distance "r = 1m" is given as follows:


"E = k\dfrac{q}{r^2}"

where "k = 9\times 10^{9}N\cdot m^2/C^2" is the Coulomb's constant. Thus, obtain:


"E = 9\times 10^{9}\dfrac{N\cdot m^2}{C^2}\cdot \dfrac{45\times 10^{-6}C}{(1m)^2} = 4.05\times 10^5N/C"

Answer. "4.05\times 10^5N/C".